# Geometry-aware 4D Video Generation for Robot Manipulation

### v1

@@ 4.3 Robot Policy Results
We evaluate the robot policy’s accuracy and generalization to novel camera views across three simulated tasks. Each task is tested 30 times with varying object positions and camera viewpoints. The success rate of task completion is reported in Table 2.
@@
The video generation model takes RGB-D observations from two novel camera views as input and predicts future observations. The generated 4D video is then passed to the pose tracking model to extract 6DoF gripper poses for both robot arms. Gripper openness is inferred using the method described in Section 3.4. The robot executes actions in an open-loop manner: after each execution, the next inference is performed using the updated RGB-D observations. Each inference takes approximately 30 seconds to generate 10 future steps on 1 NVIDIA GeForce RTX 4090 GPU.
@@
Baselines. We train baselines on the same dataset described in Section 4.1 as our method, and test on novel camera views during deployment.
@@
Dreamitate [32]: a state-of-the-art video generation method for visuomotor policy. Dreamitate uses a pretrained SVD model and finetunes it on robotic task videos to generate stereo RGB video predictions. Since it does not predict depth, it employs MegaPose [43] to extract the 6DoF pose of end effectors from the generated videos.
@@
Diffusion Policy [42]: a UNet diffusion model that predicts future robot end-effector trajectories, conditioned on history RGB image observations from two camera views. We randomly sample two camera views in the training dataset and encode their corresponding RGB observations using a CLIP-pretrained ViT model [44], following the same encoder setup as in [42]. The image features are then concatenated and fed as a condition to the U-Net diffusion model to predict future actions. The diffusion model outputs robot end effector trajectories in the next 16 steps. The model is evaluated on two unseen camera views.
@@
As shown in Table 2, Dreamitate consistently underperforms across all tasks. The lack of depth prediction and geometric consistency supervision results in lower-quality, view-inconsistent video outputs, which degrade the accuracy of the extracted poses. Consequently, the downstream robot policy suffers from a high failure rate. In contrast, our method—which jointly predicts RGB-D sequences and enforces 3D consistency—achieves over 55% higher task success rate on average.
In addition, Diffusion Policy struggles to generalize to unseen viewpoints, even though it is trained on demonstrations from multiple views. This is because the model does not explicitly model geometric correspondences across views and it’s challenging for the model to learn view-invariant actions by simply conditioning on features extracted from multi-view images.
@@ 5 Limitation
First, our method requires an RGB-D video dataset with varying camera viewpoints for training. While such datasets are easy to generate in simulation, collecting them in the real world is challenging due to hardware constraints and camera calibration requirements. Additionally, most of our experiments assume clean depth images; however, obtaining high-quality depth in real-world settings is often difficult. Recent advances in bridging the sim-to-real gap for depth data, such as FoundationStereo [45], show promising results and could be leveraged to support real-world data generation in future work. Second, the inference speed of the current video generation model is relatively slow. Recent flow matching [46, 47] or autoregressive transformers [48, 49, 50, 51] have demonstrated faster inference speed, which could lead to more reactive robot policies in future works.
@@ 6 Conclusion
We present a 4D video generation model that produces spatio-temporally consistent RGB-D sequences. Our method introduces geometric-consistent supervision during training by projecting pointmaps from one camera view into another to enforce cross-view consistency. By learning a shared geometric space, the model can generate future RGB-D videos from novel viewpoints without requiring camera poses at inference time. We demonstrate improved video generation quality and 3D consistency compared to baseline methods. Additionally, the generated 4D videos can be directly used to extract robot actions using an off-the-shelf 6DoF pose tracking model, enabling effective execution of robot manipulation tasks.
@@ Acknowledgments
Appendix A Technical Appendices and Supplementary Material
In Section A.1, we provide more details of our 4D generation model architecture. In Section A.2, we describe the camera sampling method used to generate the multi-view RGBD video dataset for training the 4D generation model. In Section A.3, we provide details such as compute resources requirements and hyperparameter choices for model training and inference. In Section A.4, we provide more quantitative and qualitative results of our method and additional baselines. In Section A.5, we discuss the broader impact of our work.
@@ A.1 Model Details
In §3.1 and §3.2, we discussed that the model takes in historical video frames $\{\mathbf{O}_{t-h+1},\cdots,\mathbf{O}_{t}\}$ and historical pointmaps $\{\mathbf{X}_{t-h+1},\cdots,\mathbf{X}_{t}\}$ from both the native view $v_{n}$ and the second view $v_{m}$. In practice, we use the latest observation, and repeat it $h=10$ times to match the number of frames that needs to be predicted, following the implementation in SVD [11].
Each pair of RGB image and pointmap condition ${\mathbf{O}^{v}_{t},\mathbf{X}^{v}_{t}}$, where $v\in\{n,m\}$, is independently encoded using separate VAE encoders for images and pointmaps, as detailed in §3. The image VAE encodes each RGB frame into a latent feature of shape $h\times c\times w^{\prime}\times h^{\prime}$, where $h=10$ is the temporal horizon, $c{=}4$ is the latent channel size, $w^{\prime}{=}32$ and $h^{\prime}{=}40$ are spatial dimensions of the latent feature maps. Similarly, the pointmap VAE encodes pointmap into shape $h\times c\times w^{\prime}\times h^{\prime}$. These encoded image and pointmap features are then concatenated along the channel axis with the corresponding noisy latents of future images and pointmaps, yielding a combined input tensor of shape $h\times 4c\times w^{\prime}\times h^{\prime}$ ($h\times 16\times 32\times 40$), which is fed into the U-Net diffusion model.
@@
To allow information sharing between the two diffusion branches as shown in Figure 2, we add one cross-attention layer after each decoder block in the U-Net diffusion model for the branch corresponding to view $v_{m}$. This results in 12 added cross attention layers. As illustrated in Figure 6, the query to the cross-attention layer are feature map tokens (feature at each pixel in the feature map) output by the decoder block in view $v_{m}$, where $c^{\prime}$ is the feature dimension, $h^{\prime}$ and $w^{\prime}$ are spatial dimensions of the feature map; the key and value are feature map tokens output by the corresponding decoder block in the native view $v_{n}$’s branch. The updated features are passed to the next decoder block in view $v_{m}$. The cross-attention layers capture spatial correspondences between the views through our geometric-consistent supervision mechanism.
We use pre-trained weights in SVD to initialize the denoising U-Net model and find that using pre-trained weights helps the model converge faster. To get better prediction quality around the robot gripper, which is important for action extraction later, we apply a re-weighting mechanism in the diffusion loss. Concretely, we use binary masks (in simulation, the object segmentations are provided; in real world, we use SAM2 [37]) of the robot gripper region and downsample it by a factor of 8 to match the resolution of the latent space while still keeping the spatial correspondence. The resulting downsampled masks provide a spatial weight map at each timestep $t^{\prime}$, denoted as $w_{g}(t^{\prime})$, which is incorporated into the joint diffusion loss mentioned in §3.3 as follows:
@@ A.2 Dataset Details
To sample cameras for rendering multi-view RGBD videos, we first sample camera positions within a half-sphere shell defined by an inner radius ($r_{1}=0.7m$) and outer radius ($r_{2}=1.2m$), with the center being the origin of the world coordinate system (center of the table). We restrict the range of the camera positions within the area between $0.2m\leq x\leq 0.6m$, $-0.5m\leq y\leq 0.5m$, and $0.7\leq z\leq 1.2m$, as shown in Figure 7 (a-c). The world coordinate system is shown in Figure 7 (d).
@@ A.3 Training Details
The 4D generation model described in § 3 is trained separately for each task in § 4 for approximately 60 epochs using 4 NVIDIA RTX A6000 GPUs (48GB memory each). We fine-tune the full U-Net backbone of the SVD [11] model with a learning rate of $1\times 10^{-6}$, using the AdamW optimizer ($\beta_{1}=0.95$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, weight decay $=1\times 10^{-6}$) and a batch size of 4. The image and pointmap VAE encoders are frozen during diffusion model training.
At inference time, we apply the standard EulerEDMSampler [52] with 25 denoising steps. For robot policy deployment, both the generation model and the pose tracking model are run on a single NVIDIA GeForce RTX 4090 GPU.
@@ A.4 Additional Results
Figure 8: Qualitative Multi-View Video Generation Results. We show temporal results generated by our 4D video generation model across three robot manipulation tasks. With geometry-consistent supervision and joint temporal and 3D consistency optimization, our model is able to output spatio-temporally consistent videos across camera views with high visual fidelity.
@@
In Figure 8, we show generated RGB video sequences for the StoreCerealBoxUnderShelf, PutSpatulaOnTable, and PlaceAppleFromBowlIntoBin tasks using our proposed 4D generation model. With geometry-consistent supervision and joint temporal and 3D consistency optimization, our model is able to output spatio-temporally consistent videos across camera views with high visual fidelity. We also show baseline comparison results on the PlaceAppleFromBowlIntoBin task in Figure 9 and PutSpatulaOnTable task in Figure 10. Our method consistently achieves the best RGB video and depth generation quality, with high multi-view consistency. Baseline results often exhibit significant cross-view inconsistencies (marked in red) or contain noticeable artifacts in the RGB or depth predictions.
A.5 Broader Impact
Positive impacts. Our 4D video generation model can improve robot manipulation by providing consistent, multi-view RGB-D predictions for pose tracking, and provide better interpretability comparing to typical behavior cloning approaches. The model can be used to enhance the capabilities of general-purpose robotic systems in household, factories, etc.
Negative impacts. Like other generative models, our method could be misused for creating realistic but deceptive content (e.g., deepfakes). If used in critical settings without robust validation, errors in spatial prediction could lead to unsafe robot behaviors. Dataset biases may also limit generalization.
Mitigations. We recommend controlled model release, output watermarking, and bias-aware dataset design. In robotics, deployment should include safety checks and fallback mechanisms. Future work should explore interpretability and robustness under real-world conditions.

### v2

@@ 4.3 Robot Policy Results
We evaluate the robot policy’s success rate and generalization to novel camera views across three simulation tasks. Each task is tested in 30 rollouts with unseen initial object poses and novel camera viewpoints sampled from the test dataset described in Section˜4.1. The success rate of task completion is reported in Table˜2.
@@
The video generation model takes RGB-D observations from two novel camera views as input and predicts future observations. The generated 4D video is then passed to the pose tracking model to extract 6DoF gripper poses for both robot arms. Gripper openness is inferred using the method described in Section˜3.4. The robot executes actions in an open-loop manner: after each execution, the next inference is performed using the updated RGB-D observations. Each inference takes approximately 30 seconds to generate 10 future frames on 1 NVIDIA GeForce RTX 4090 GPU.
@@
Baselines. We train baselines on the same dataset described in Section˜4.1 as our method, and test on novel camera views during deployment.
@@
Dreamitate (Liang et al., 2024): a state-of-the-art video generation method for visuomotor policy. Dreamitate uses a pretrained SVD model and finetunes it on robotic task videos to generate stereo RGB video predictions. Since it does not predict depth, it employs MegaPose (Labbé et al., 2022) to extract the 6DoF pose of end effectors from the generated videos.
@@
Diffusion Policy (DP) (Chi et al., 2023): a UNet diffusion model that predicts future robot end-effector trajectories, conditioned on history RGB image observations from two camera views. We randomly sample two camera views in the training dataset and encode their corresponding RGB observations using a CLIP-pretrained ViT model (Radford et al., 2021), following the same encoder setup as in (Chi et al., 2023). The diffusion model outputs robot end effector trajectories in the next 16 steps of which the first 6 actions are executed. The model is evaluated on two unseen camera views sampled from the test dataset.
DP3 (Ze et al., 2024): To compare our method with a behavior cloning method that takes in RGB-D information, we implement DP3, a UNet-based diffusion model that predicts future robot end-effector trajectories conditioned on global 3D point clouds with colors. Following the DP baseline setup, we randomly sample RGB-D images from two training camera views and aggregate their projected global point clouds as input condition. The evaluation setting is same as the DP baseline.
@@
As shown in Table˜2, Dreamitate consistently underperforms across all tasks. The lack of depth prediction and geometric consistency supervision results in lower-quality, view-inconsistent video outputs, which degrade the accuracy of the extracted poses. Consequently, the downstream robot policy suffers from a high failure rate. In contrast, our method—which jointly predicts RGB-D sequences and enforces 3D consistency—achieves over 50% higher task success rate on average.
In addition, Diffusion Policy struggles to generalize to unseen viewpoints, even though it is trained on demonstrations from multiple views. This is because the model does not explicitly model geometric correspondences across views and it’s challenging for the model to learn view-invariant actions by simply conditioning on features extracted from multi-view images. Using global point clouds as input, DP3 achieves slightly better performance on the StoreCerealBoxUnderShelf task, improving cereal box grasp accuracy with the added depth information. However, no gains are observed on the other two tasks, likely because the objects being manipulated are too small and the benefit of end-to-end learning from point clouds is not obvious. Most failures remain grasping failures on the spatula or apple. On the other hand, direct gripper pose tracking on generated RGB-D videos yields higher action accuracy and improves grasping success.
@@ 5 Limitation
First, our method requires RGB-D video dataset with varying camera viewpoints for training. While such datasets are easy to generate in simulation, collecting them in the real world is challenging due to hardware constraints and camera calibration requirements. Additionally, obtaining high-quality depth in real-world settings is often difficult. Recent advances in depth estimation from RGB images (Wen et al., 2025; Wang et al., 2025) show promising results and can be leveraged to support high quality data curation in real-world environments in future work. Second, the inference speed of the current video generation model is relatively slow compared to an end-to-end behavior cloning policy, making closed-loop planning difficult. Recent flow matching (Davtyan et al., 2023; Jin et al., 2024) or autoregressive transformers (Yin et al., 2024; Deng et al., 2024; Li et al., 2024b; Gu et al., 2025) have demonstrated faster video generation speed, which could lead to more reactive robot policies.
@@ 6 Conclusion
We present a 4D video generation model that produces spatio-temporally consistent RGB-D sequences. Our method introduces geometric-consistent supervision during training by projecting pointmaps from one camera view into another to enforce cross-view consistency. By learning a shared geometric space, the model can generate future RGB-D videos from novel viewpoints without requiring camera poses at inference time. We demonstrate improved video generation quality and 3D consistency compared to baseline methods. Additionally, the generated 4D videos can be directly used to extract robot actions using an off-the-shelf 6DoF pose tracking model, enabling higher success rate on several robot manipulation tasks.
7 Ethics statement
This work focuses on developing models for multi-view video generation and their application to robot manipulation. The primary goal is to advance general-purpose robotic capabilities in household and assistive settings. The datasets used in this work include synthetic data generated in simulation and real-world robot demonstrations collected in controlled lab environments; no human subjects or sensitive personal data are involved. The proposed method does not generate or rely on human biometric information. Potential societal benefits include enabling robots to better assist with daily tasks, elder care, and other beneficial applications. At the same time, we acknowledge potential misuse risks, such as manipulative content generation. To mitigate these risks, we limit the scope of this work to robotic data and will release code and models under a license that restricts harmful use.
8 Reproducibility statement
We have made extensive efforts to ensure the reproducibility of our work. Details of the model architecture, training objectives, and hyperparameters are provided in Section˜3 in the main paper and Section˜A.1, Section˜A.3, Section˜A.4 in Appendix. The datasets used in simulation and real-world experiments, including task design and camera sampling procedure, are described in Appendix Section˜A.2. Evaluation metrics and baseline implementations are specified in Section˜4 in the main paper. We also provide a project website (https://robot4dgen.github.io/) with qualitative results and visualizations, and we will release code, datasets, and pre-trained models upon acceptance to further support reproducibility.
@@ Acknowledgments
Appendix A Appendix
In Section˜A.1, we provide more details of our 4D generation model architecture. In Section˜A.2, we describe the camera sampling method used to generate the multi-view RGBD video dataset for training the 4D generation model. In Section˜A.3, we provide details such as compute resources requirements and hyperparameter choices for model training and inference. In Section˜A.4, we provide more quantitative and qualitative results of our method and additional baselines. In Section˜A.5, we discuss the LLM usage of our work.
@@ A.1 Model Details
In §3.1 and §3.2, we discussed that the model takes in historical video frames $\{\mathbf{O}_{t-h+1},\cdots,\mathbf{O}_{t}\}$ and historical pointmaps $\{\mathbf{X}_{t-h+1},\cdots,\mathbf{X}_{t}\}$ from both the native view $v_{n}$ and the second view $v_{m}$. In practice, we use the latest observation, and repeat it $h=10$ times to match the number of frames that needs to be predicted, following the implementation in SVD (Blattmann et al., 2023).
Each pair of RGB image and pointmap condition $\mathbf{O}^{v}_{t},\mathbf{X}^{v}_{t}$}, where $v\in\{n,m\}$, is independently encoded using separate VAE encoders for images and pointmaps, as detailed in §3. The image VAE encodes each RGB frame into a latent feature of shape $h\times c\times w^{\prime}\times h^{\prime}$, where $h=10$ is the temporal horizon, c=4c{=}4 is the latent channel size, $w^{\prime}{=}32$ and $h^{\prime}{=}40$ are spatial dimensions of the latent feature maps. Similarly, the pointmap VAE encodes pointmap into shape $h\times c\times w^{\prime}\times h^{\prime}$. These encoded image and pointmap features are then concatenated along the channel axis with the corresponding noisy latents of future images and pointmaps, yielding a combined input tensor of shape $h\times 4c\times w^{\prime}\times h^{\prime}$ ($h\times 16\times 32\times 40$), which is fed into the U-Net diffusion model.
@@
To allow information sharing between the two diffusion branches as shown in Figure˜2, we add one cross-attention layer after each decoder block in the U-Net diffusion model for the branch corresponding to view $v_{m}$. This results in 12 added cross attention layers. As illustrated in Figure˜5, the query to the cross-attention layer are feature map tokens (feature at each pixel in the feature map) output by the decoder block in view $v_{m}$, where $c^{\prime}$ is the feature dimension, $h^{\prime}$ and $w^{\prime}$ are spatial dimensions of the feature map; the key and value are feature map tokens output by the corresponding decoder block in the native view $v_{n}$’s branch. The updated features are passed to the next decoder block in view $v_{m}$. The cross-attention layers capture spatial correspondences between the views through our geometric-consistent supervision mechanism.
We use pre-trained weights in SVD to initialize the denoising U-Net model and find that using pre-trained weights helps the model converge faster. To get better prediction quality around the robot gripper, which is important for action extraction later, we apply a re-weighting mechanism in the diffusion loss. Concretely, we use binary masks (in simulation, the object segmentations are provided; in real world, we use SAM2 (Ravi et al., 2024)) of the robot gripper region and downsample it by a factor of 8 to match the resolution of the latent space while still keeping the spatial correspondence. The resulting downsampled masks provide a spatial weight map at each timestep $t^{\prime}$, denoted as $w_{g}(t^{\prime})$, which is incorporated into the joint diffusion loss mentioned in §3.3 as follows:
@@ A.2 Dataset Details
A.2.1 Simulation Tasks
Figure 6: Simulation Tasks for Evaluation.
A.2.2 Camera Sampling
To sample cameras for rendering multi-view RGBD videos, we first sample camera positions within a half-sphere shell defined by an inner radius ($r_{1}=0.7m$) and outer radius ($r_{2}=1.2m$), with the center being the origin of the world coordinate system (center of the table). We restrict the range of the camera positions within the area between $0.2m\leq x\leq 0.6m$, $0.5m\leq y\leq 0.5m$, and $0.7\leq z\leq 1.2m$, as shown in Figure˜7 (a-c). The world coordinate system is shown in Figure˜7 (d).
@@ A.2.2 Camera Sampling
A.2.3 Real World Tasks
Figure 8: Tasks in Real-World Dataset.
@@ A.3 Training Details
The 4D generation model described in § 3 is trained separately for each task in § 4 for approximately 60 epochs using 4 NVIDIA RTX A6000 GPUs (48GB memory each). We fine-tune the full U-Net backbone of the SVD (Blattmann et al., 2023) model with a learning rate of $1\times 10^{-6}$, using the AdamW optimizer ($\beta_{1}=0.95$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, weight decay $1\times 10^{-6}$) and a batch size of 4. The image and pointmap VAE encoders are frozen during diffusion model training.
At inference time, we apply the standard EulerEDMSampler (Karras et al., 2022) with 25 denoising steps. For robot policy deployment, both the generation model and the pose tracking model are run on a single NVIDIA GeForce RTX 4090 GPU.
Method Inference Time Training Memory # of Trainable Parameters
OURS 30.0 s 47 G 2.4 B
OURS w/o MV attn 29.3 s 46.5 G 2.38 B
4D Gaussian 2 s 2813 M 856,774
SVD (Dreamitate) 13.4 s 45.8 G 1.54 B
SVD w/ MV attn 15.1 s 46.3 G 2.4 B
Table 3: Quantitative comparisons with baselines in terms of inference time, GPU memory / compute resource consumption and parameter count.
We report quantitative comparisons with baselines in terms of inference latency, GPU memory / compute resource consumption, and parameter count. Our method requires 30.0s inference time with 47G training memory and 2.4B parameters: adding multi-view cross attention layers only slightly increases each metric yet yields more geometry-consistent and high-quality generation results. In contrast, 4D Gaussian is significantly faster (2s) and lighter (2813M memory, 856,774 parameters), but produces poor quality results for novel view rendering and also requires a generated single-view video as input, while our method directly predicts future frames based on a single RGB-D image per view.
We acknowledge that the higher inference time reflects the cost of predicting both RGB and depth modalities, but this trade-off results in stronger video fidelity and geometric consistency. Looking ahead, improvements in efficiency can be explored by applying the geometry consistency objective to faster generative models, such as leveraging 3D VAEs (Zhao et al., 2024) to reduce temporal dimensions of the letent tokens, adopting one-step or few-step diffusion models (Lin et al., 2025), or employing autoregressive transformers with lower latency (Huang et al., 2025b). On the policy side, a hierarchical structure (Black et al., 2023; Wen et al., 2023) could further mitigate the cost of dense video inference: instead of extracting and executing all poses from generated RGB-D frames, the model can act as a high-level planner where sparse predicted frames serve as subgoals, while a low-level controller or learned policy fills in reactive actions. This design would reduce the number of generated frames required for action planning, thereby accelerating inference while still preserving the advantages of geometry-aware video generation for downstream robotic tasks.
@@ A.4 Additional Results
Figure 9: Qualitative Multi-View Video Generation Results in Simulation. We show temporal results generated by our 4D video generation model across three robot manipulation tasks. With geometry-consistent supervision and joint temporal and 3D consistency optimization, our model is able to output spatio-temporally consistent videos across camera views with high visual fidelity in unseen views.
Figure 10: Qualitative Multi-View Video Generation Results in Real World. We show temporal results generated by our 4D video generation model across four real-world robot manipulation tasks. By fine-tuning the simulation-trained model for around 15k steps, our model is able to output spatio-temporally consistent videos across camera views with high visual fidelity.
@@
In Figure˜9 and Figure˜10, we show generated RGB video sequences on both simulation and real world tasks using our proposed 4D generation model. With geometry-consistent supervision and joint temporal and 3D consistency optimization, our model is able to output spatio-temporally consistent videos across camera views with high visual fidelity. We also show baseline comparison results on the PlaceAppleFromBowlIntoBin task in Figure˜11 and PutSpatulaOnTable task in Figure˜12. Our method consistently achieves the best RGB video and depth generation quality, with high multi-view consistency. Baseline results often exhibit significant cross-view inconsistencies (marked in red) or contain noticeable artifacts in the RGB or depth predictions.
A.5 The Use of Large Language Models (LLMs)
We used large language models (LLMs), specifically OpenAI’s ChatGPT 5, to assist with writing and editing the manuscript. The models were used primarily for improving the clarity, grammar, and readability of the text, including rephrasing sentences, polishing explanations, and standardizing style. Anthropic’s Claude was used to assist in generating visualization-related code, such as for camera pose and depthmap visualization in the figures. All technical content, experimental design, analysis, and conclusions were conceived, implemented, and validated by the authors without the use of LLMs. The authors have carefully reviewed and verified all content produced with the assistance of LLMs to ensure accuracy and originality.
A.6 Pose Tracking Error Analysis
The pose tracking error arises from two sources: (1) the RGB-D generation quality of the video model, and (2) the detection accuracy of FoundationPose. To disentangle these factors, we additionally evaluate FoundationPose directly on ground-truth RGB-D images to isolate its performance independently of any generation errors. As shown in Figure˜13, we visualize one representative trajectory from each task and report the per-action MSE between the detected pose and the ground-truth pose throughout the task. We also include the frames with the top five highest pose-tracking errors to illustrate the characteristic failure cases.
We observe that the MSE error of pose tracking is generally around or below 0.001 for both grippers across tasks, indicating good pose-tracking accuracy. The only notable exception is the right gripper in the PutSpatulaOnTable task and the left gripper in StoreCerealBoxUnderShelf towards the end of the task, where the error occasionally peaks at around 0.003.
Figure 13: Analysis of 6-DoF pose-tracking errors. Left: per-action MSE curves for left and right grippers across one representative trajectory from each of the three tasks. Right: visual examples of the five highest-error frames for each task.
A.7 Scaling to More Input Views
Our main evaluation focuses on simultaneously generating two views. We additionally study how the proposed framework scales when incorporating more input views at inference time.
Without retraining, the model can incorporate a third (or more) view(s) at inference time by performing an additional forward pass between the reference view and each new view. This approach increases inference time approximately linearly with the number of views, but does not require architectural modifications or additional training. Designing more efficient mechanisms for large-scale multi-view inference is an interesting direction for future work, similar in spirit to how Fast3R (Yang et al., 2025a) extends DUSt3R to handle thousands of input images.
View mIoU FVD AbsRel $\bm{\delta_{1}}$
View 1 515.53 0.06 0.95
View 2 (w/ View 1) 0.62 566.86 0.09 0.95
View 3 (w/ View 1) 0.54 662.83 0.12 0.89
Table 4: Performance generating three views simultaneously on StoreCerealBoxUnderShelf.
Three-view Evaluation. We evaluated the performance of incorporating a third predicted view on the StoreCerealBoxUnderShelf task. In this setting, the model simultaneously generates videos in three randomly sampled, unseen viewpoints, and the reported metrics are averaged across different initial RGB-D conditions.
Table 4 summarizes the results. Although performance degrades slightly, the model is still able to produce RGB-D videos with good visual quality and cross-view consistency.